A kettle transfers 6000 J of energy electrically. 1500 J of this is wasted. What is the efficiency of this kettle?
1
Expert's answer
2020-06-02T17:53:05-0400

We can find the efficiency of the kettle from the formula:


"Efficiency = \dfrac{useful \ power \ transferred}{total \ power \ supplied} \cdot 100\%."

The useful power transferred by the kettle can be found from the formula:


"Useful \ power \ transferred = Supplied \ electrical \ energy - wasted \ energy,""Useful \ power \ transferred = 6000J - 1500J = 4500J."

Then, we can calculate the efficiency of the kettle:


"Efficiency = \dfrac{4500J}{6000J} \cdot 100\% = 75\%."

Answer:

"Efficiency = 75\%."